Presented below are selected transactions for Bridgeport Company during September and October of the current year. Bridgeport uses a periodic inventory system.

Sept. 1 Purchased merchandise on account from Hillary Company at a cost of $45,000, FOB destination, terms 1/15, n/30.
2 The correct company paid $2,000 of freight charges to Trucking Company on the September 1 merchandise purchase.
5 Returned for credit $3,000 of damaged goods purchased from Hillary Company on September 1.
15 Sold the remaining merchandise purchased from Hillary Company to Irvine Company for $82,600, terms 2/10, n/30, FOB destination.
16 The correct company paid $2,200 of freight charges on the September 15 sale of merchandise.
17 Issued Irvine Company a credit of $5,900 for returned goods. These goods had cost Bridgeport Company $3,000 and were returned to inventory.
25 Received the balance owing from Irvine Company for the September 15 sale.
30 Paid Hillary Company the balance owing for the September 1 purchase.
Oct. 1 Purchased merchandise on account from Kimmel Company at a cost of $52,000, terms 2/10, n/30, FOB shipping point.
2 The correct company paid freight costs of $1,100 on the October 1 purchase.
3 Obtained a purchase allowance of $2,000 from Kimmel Company to compensate for some minor damage to goods purchased on October 1.
10 Paid Kimmel Company the amount owing on the October 1 purchase.
11 Sold all of the merchandise purchased from Kimmel Company to Kieso Company for $98,500, terms 2/10, n/30, FOB shipping point.
12 The correct company paid $800 freight costs on the October 11 sale.
17 Issued Kieso Company a sales allowance of $2,000 because some of the goods did not meet Kieso's exact specifications.
31 Received a cheque from Kieso Company for the balance owing on the October 11 sale.



Record the September and October transactions for Bridgeport Company. Assume that Norlan uses the earnings approach. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Record journal entries in the order presented in the problem.)

Date

Account Titles and Explanation

Debit

Credit
Answer #1

Answer:

Date Account Titles Debit Credit
Sep-01 Inventory $            45000
         Accounts Payable $          45000
Sep-02 no entry
Sep-05 Accounts Payable $               3000
       Inventory $            3000
Sep-15 Accounts Receivable $            82600
       Sales Revenue $          82600
Cost of Goods Sold $            42000 =45000-3000
      Inventory $          42000
Sep-16 Delivery Expense $               2200
      Cash $            2200
Sep-17 Sales returns and allowances $               5900
       Accounts Receivable $            5900
Inventory $               3000
      Cost of Goods Sold $            3000
Sep-25 Cash $            75166 =76700-1534
Sales Discount $               1534 =76700*2%
       Accounts Receivable $          76700
Sep-30 Accounts Payable $            42000 =45000-3000
       Cash $          42000
Oct-01 Inventory $            52000
      Accounts Payable $          52000
Oct-02 Inventory $               1100
      Cash $            1100
Oct-03 Accounts Payable $               2000
      Inventory $            2000
Oct-10 Accounts Payable $            50000 =52000-2000
      Cash $          49000 =50000-1000
      Inventory $            1000 =50000*2%
Oct-11 Accounts Receivable $            98500
       Sales Revenue $          98500
Cost of Goods Sold $            50100 =52000+1100-2000-1000
      Inventory $          50100
Oct-12 no entry
Oct-17 Sales Returns and allowances $               2000
     Accounts Receivable $            2000
Oct-31 Cash $            96500 =98500-2000
     Accounts Receivable $          96500